# Clustering Algorithms and RAG Enhancing Semi-Supervised Text Classification with Large LLMs

### v1

@@ 6 Experiment
Figure 4 further illustrates the impact of text augmentation on Jaccard Similarity (vocabulary overlap) among generated samples. From RAG augmentation to LLMs rewrite to traditional WordNet random synonym replacement, the vocabulary overlap decreases, indicating the generation of more diverse samples. When comparing the Jaccard Similarity between generated samples and the original documents, RAG augmentation achieved the lowest scores, averaging 42.31% for the Reuters dataset and 36.08% for the WOS dataset. Despite the lower vocabulary overlap, the embedding-based similarity remained high, with scores of 92.17% for the Reuters dataset and 90.84% for the WOS dataset. This suggests that while RAG augmentation introduces significant changes, it maintains the core meaning and semantic coherence of the original text. The RAG method stands out for its ability to introduce new information while preserving the original context, making it ideal for data augmentation in machine learning and content diversification in NLP tasks. The high embedding similarity scores further validate its effectiveness in maintaining semantic integrity despite significant textual modifications.
@@ 7.1 Landmark vs Random
  • Domain Match: The generated domain matched with the true domain
  • Area Match: The generated area matched with the true area
To evaluate the effectiveness of our clustering techniques compared to random selection, we conducted a comparative analysis using two main types of data: landmark data and random data. The landmark data was carefully selected to represent meaningful and representative samples from the dataset, reflecting core themes. In contrast, the random data was chosen without specific criteria, serving as a baseline to gauge the added value of the clustering technique. To ensure a fair assessment of randomness, we performed three trials for the random data due to computational constraints. The datasets used in this study are detailed below:
  • Landmark Data: Selected landmarks believed to be indicative of the core themes within each dataset, 300 from Reuters and 600 from WoS. These samples were trained for 15 epochs, which yielded the highest performance.
  • Random Data: An equivalent number of samples (300 from Reuters and 600 from WoS) randomly chosen from the training set to serve as a control group. They were also trained for 15 epochs, achieving optimal performance.
  • Landmark-WordNet Data: The landmark data was augmented tenfold using WordNet, expanding the dataset to 3,000 samples from Reuters and 6,000 from WoS. Train for 1.5 epoch.
  • Random-WordNet Data: Similarly, the random data was also augmented tenfold using WordNet, resulting in 3,000 samples from Reuters and 6,000 from WoS. Train for 1.5 epoch.
Dataset Part Match All Match In Right Order
Trail 1, random data 88.92% 80.81% 74.28%
Trail 2, random data 88.13% 82.90% 76.08%
Trail 3, random data 87.85% 80.76% 74.23%
Landmark data 90.85% 82.35% 75.86%
Trail 1, random wordnet data 89.52% 82.25% 76.43%
Trail 2, random wordnet data 89.57% 83.45% 75.57%
Trail 3, random wordnet data 88.13% 81.74% 73.54%
Landmark wordnet data 93.05% 83.97% 78.82%
The results, as summarized in Table 7 for the Reuters dataset, reveal several interesting insights. Initially, it might seem counterintuitive that in some trials, random data outperformed landmark data in terms of "All match" and "In right order" accuracy. However, this phenomenon can be attributed to the fact that when the data volume is small, more data about the majority classes is needed to improve model performance. During random selection, more samples are likely to fall into the majority class. Additionally, the test set contains more data representing common cases rather than edge cases. Therefore, in the initial stage of improving accuracy, more data describing the majority of the dataset is required to ensure the model generalizes well to typical scenarios. Despite this initial finding, the true value of the landmark data becomes evident when the datasets are augmented with WordNet. The simple augmentation process generated more samples for each category, which in turn significantly boosted the performance of the landmark data over the random data.
For instance, the Part Match metric for the Reuters landmark data improved from 90.85% to 93.05% after augmentation, representing a 2.2% enhancement in the model’s ability to generate at least one relevant tag. This is significantly higher than the 0.77% average improvement seen with random data. Similarly, the All Match and In Right Order metrics also showed substantial gains, increasing from 82.35% to 83.97% (1.62%) and from 75.86% to 78.82% (2.95%), respectively. In contrast, the average improvements for All Match and In Right Order with random data were only 0.99% and 0.31%. These results indicate that the clustering technique used to select the landmark data provides more meaningful and valuable insights for enhancing model performance. While random data can sometimes offer better initial performance, the strategic use of landmark data and semantic augmentation through resources like WordNet can lead to more robust and accurate tag generation.
Domain Match Area Match
Trail 1, random data 74.84% 47.13%
Trail 2, random data 77.25% 47.09%
Trail 3, random data 72.03% 46.10%
Landmark data 73.85% 47.48%
Trail 1, random wordnet data 76.99% 51.21%
Trail 2, random wordnet data 77.32% 50.57%
Trail 3, random wordnet data 77.30% 50.06%
Landmark wordnet data 77.35% 52.09%
Table 8: Performance Comparison: Landmark vs. Random Data for WOS.
Similarily, for the results shown in Table 8, the landmark-WordNet data achieved the highest accuracy in both Domain Match (77.35%) and Area Match (52.09%), with 3.5% and 4.61% increase. Without augmentation, the landmark data slightly underperformed compared to random data in both metrics. Similar to the reuter dataset, this could be because the landmark data selection focuses on capturing core themes within the dataset, which may not always align perfectly with the overall distribution of the majority of samples. In contrast, random data selection tends to better reflect the overall distribution of the dataset, especially when the sample size is small, making it more likely to cover the main categories. After applying WordNet augmentation, there is a significant performance improvement for both landmark and random data in both Domain Match and Area Match. Notably, the landmark-WordNet data shows a more pronounced enhancement compared to the random-WordNet data. The significant improvement in Area Match for the landmark-WordNet data is particularly noteworthy as area Match is a more granular metric that requires higher precision from the model. While unenhanced landmark data may sometimes perform less well than random data, the value of landmark data becomes evident when augmented with semantic resources like WordNet and the LLMs rewrite and RAG augmentation methods we later shown.
7.2 LLMs Rewrite and RAG augmentation
@@
We then investigate the performance of large language models (LLMs) for rewriting and RAG-augmented data. This investigation aims to understand how these methods can enhance the quality and diversity of the training data, potentially leading to further improvements in model performance.
@@
  • CoT+RAG: This method involves using a larger Qwen 2.5 72B model with RAG to provide more information and CoT prompt instructions to sequentially generate answers. The prompts used for this process are detailed in Table 3.3, and no parameter tuning was involved.
@@
  • Rewrite Data: In this approach, the Landmark Data is rewritten using a Qwen 2.5 72B language model, as detailed in Table 3.2. This expands the dataset to 3,000 samples from Reuters and 6,000 from Web of Science (WoS).
@@
  • RAG Augmented Data: In this approach, each document is augmented using a Qwen 2.5 72B language model, as detailed in Table 3.4. This expands the dataset to 15,473 samples from Reuters and 62,886 from WoS.
@@
  • Combined Data: This method combines the WordNet, LLM rewriting, and RAG augmentation techniques by merging the training data together.
@@
From the results presented in Table 10, we observe a less pronounced performance improvement for RAG augmentation compared to the Reuters dataset. Specifically, even with the RAG Augmented Data method, the performance in terms of Domain Match and Area Match only marginally exceeds that of the Landmark WordNet data. This can be attributed to several factors: research topics are inherently more complex and specialized than news articles, making the LLM rewriting and RAG generation processes less effective compared to traditional WordNet methods, which typically involve replacing only certain "safe" words. Additionally, the terminology and concepts found in scientific literature are more technical and abstract, posing greater challenges for the model to accurately understand and generate high-quality text.
@@
Despite these challenges, fine-tuned models consistently outperform the RAG-CoT approach without parameter tuning, highlights the importance of task-specific model fine-tuning. Moreover, it is clear that using all original training data, which includes over 20,000 labeled samples, leads to significantly better overall performance. However, in real-world scenarios where labeled data is often limited, our methods for generating additional augmented data still contribute to huge boost in performance improvements. This emphasizes the significance of data augmentation techniques in enhancing model performance, especially when sample sizes are constrained. Notably, combining the Landmark, rewrite, and RAG Augment data results in a substantial improvement, with an 8.58% increase in Domain Match and a 16.33% increase in Area Match compared to using only the landmark data. The significant improvement in area classification is particularly encouraging, as it suggests that the different levels of vocabulary overlap in the combined augmented data cover a broader range of case types, thereby leading to a significant performance boost when used together rather than individually. These findings not only emphasize the critical role of data augmentation in boosting model performance but also suggest that different strategies should be adopted for optimizing model performance across various types of text data.

### v3

@@ 5 Experiment
From RAG augmentation to LLMs rewrite to traditional WordNet random synonym replacement, the vocabulary overlap decreases, indicating the generation of more diverse samples. Despite this lower vocabulary overlap, the embedding-based similarity remained high, with scores of 92.17% for the Reuters dataset and 90.84% for the WOS dataset. This suggests that while RAG augmentation introduces significant changes, it maintains the core meaning and semantic coherence of the original text. For LLMs rewrites and WordNet augmentation, they reinforce the representation of the majority cases, ensuring that the model performs robustly on common scenarios without being overly influenced by edge cases. We will further demonstrate that combining documents with varying similarity levels can significantly boost prediction accuracy. This strategy leverages the core information and semantic coherence preserved in high-similarity documents while adding new perspectives from lower-similarity ones. As a result, it enhances the model’s learning and generalization, leading to more accurate predictions.
We then discuss in detail how labels were generated alongside RAG augmentation. For the Reuters dataset, we configured the LLMs to produce multiple labels for each document, with individual labels delineated by commas. In contrast, for the WoS dataset, the LLM was instructed to generate a hierarchical labeling structure, beginning with the domain and then specifying the sub-domain, with each level separated by commas. We also observed that there are cases where RAG does not perform well, such as with the "lei" category in the Reuters dataset. The "lei" category, representing the Leading Economic Index, saw no augmentation because the LLMs failed to recognize "lei" without sufficient context. A more detailed report about the distribution of the generated labels is provided in the Appendix B.
@@ 6 LLM Finetunning
  • Domain Match: The generated domain matched with the true domain.
  • Area Match: In addition to the domain matching, the generated area also matches the true area.
@@
We then investigate the performance of for these metrics under LLMs trained with different seperated datasets. This investigation aims to understand how these methods can enhance the quality and diversity of the training data, potentially leading to further improvements in model performance. The datasets used in this study are detailed below:
  • Landmark Data: Selected landmarks believed to be indicative of the core themes within each dataset, 300 from Reuters and 600 from WoS.
  • Landmark-WordNet Data: The landmark data was augmented tenfold using WordNet, expanding the dataset to 3,000 samples from Reuters and 6,000 from WoS.
@@
  • CoT+RAG: This method involves using a larger Qwen 2.5 72B model with RAG to provide more information and CoT prompt instructions to sequentially generate answers. The prompts used for this process are detailed in Appendix C.6, and no parameter tuning was involved.
@@
  • Rewrite Data: In this approach, the Landmark Data is rewritten using a Qwen 2.5 72B language model, as detailed in Appendix C.3. This expands the dataset to 3,000 samples from Reuters and 6,000 from Web of Science (WoS).
@@
  • RAG Augmented Data: In this approach, each document is augmented using a Qwen 2.5 72B language model, as detailed in Appendix C.4. This expands the dataset to 15,473 samples from Reuters and 62,886 from WoS.
@@
  • Combined Data: This method combines the WordNet, LLM rewriting, and RAG augmentation techniques by merging their training data together.
@@
From the results presented in Table 4 for the WoS dataset, we observe a less pronounced performance improvement for RAG augmentation compared to the Reuters dataset. Specifically, even with the RAG Augmented Data method, the performance in terms of Domain Match and Area Match only marginally exceeds that of the Landmark WordNet data. This aligns with our clustering results, which show that the Homogeneity score is lower when the cluster numbers are the same, and for the Silhouette Score to reach its peak, a higher percentage of data is required. This can be attributed to several factors: research topics are inherently more complex and specialized than news articles, which may render the LLM rewriting and RAG generation processes less effective compared to traditional WordNet methods. Traditional methods typically involve replacing only certain "safe" words, which are more straightforward and less error-prone. Despite these challenges, fine-tuned models consistently outperform the RAG-CoT approach without parameter tuning, highlights the importance of task-specific model fine-tuning. Moreover, it is clear that using all original training data, which includes over 20,000 labeled samples, leads to significantly better overall performance. However, in real-world scenarios where labeled data is often limited, our methods for generating additional augmented data still contribute to huge boost in performance improvements. This emphasizes the significance of data augmentation techniques in enhancing model performance, especially when sample sizes are constrained. Notably, combining the Landmark, rewrite, and RAG Augment data results in a substantial improvement, with an 8.58% increase in Domain Match and a 16.33% increase in Area Match compared to using only the landmark data. The significant improvement in area classification is particularly encouraging, as it suggests that the different levels of vocabulary overlap in the combined augmented data cover a broader range of case types, thereby leading to a significant performance boost when used together rather than individually. These findings not only emphasize the critical role of data augmentation in boosting model performance but also suggest that different strategies should be adopted for optimizing model performance across various types of text data.
7 Ablation: Random compared with Landmark data for LLMs
In addition to checking the effectiveness of the augmentation method, we evaluated our clustering techniques’ performance compared to random selection through a comparative analysis. This analysis involved two main types of data: landmark data and random data. The landmark data were carefully selected to represent meaningful and representative samples from the dataset, reflecting core themes. In contrast, the random data were chosen without specific criteria, serving as a baseline to gauge the added value of the clustering technique. To ensure a fair assessment of randomness, we conducted three trials for the random data due to computational constraints:
  • Random Data: An equivalent number of samples (300 from Reuters and 600 from WoS) randomly chosen from the training set to serve as a control group.
  • Random-WordNet Data: Similarly, the random data was also augmented tenfold using WordNet, resulting in 3,000 samples from Reuters and 6,000 from WoS.
Dataset Part Match All Match In Right Order
Trail 1, random data 88.92% 80.81% 74.28%
Trail 2, random data 88.13% 82.90% 76.08%
Trail 3, random data 87.85% 80.76% 74.23%
Landmark data 90.85% 82.35% 75.86%
Trail 1, random wordnet data 89.52% 82.25% 76.43%
Trail 2, random wordnet data 89.57% 83.45% 75.57%
Trail 3, random wordnet data 88.13% 81.74% 73.54%
Landmark wordnet data 93.05% 83.97% 78.82%
The results, as summarized in Table 5 for the Reuters dataset, reveal several interesting insights. Initially, it might seem counterintuitive that in some trials, random data outperformed landmark data in terms of "All match" and "In right order" accuracy. However, this phenomenon can be attributed to the fact that when the data volume is small, more data about the majority classes is needed to improve model performance. During random selection, more samples are likely to fall into the majority class. Additionally, the test set contains more data representing common cases rather than edge cases. Therefore, in the initial stage of improving accuracy, more data describing the majority of the dataset is required to ensure the model generalizes well to typical scenarios. Despite this initial finding, the true value of the landmark data becomes evident when the datasets are augmented with WordNet. The simple augmentation process generated more samples for each category, which in turn significantly boosted the performance of the landmark data over the random data. After augmentation, the Part Match metric for Reuters landmark data rose from 90.85% to 93.05%, a 2.2% improvement compared to a 0.77% boost with random data. Similarly, All Match increased by 1.62% (from 82.35% to 83.97%), and In Right Order by 2.95% (from 75.86% to 78.82%), significantly outpacing random data improvements of 0.99% and 0.31%, respectively. These results indicate that the clustering technique used to select the landmark data provides more meaningful and valuable insights for enhancing model performance.
Domain Match Area Match
Trail 1, random data 74.84% 47.13%
Trail 2, random data 77.25% 47.09%
Trail 3, random data 72.03% 46.10%
Landmark data 73.85% 47.48%
Trail 1, random wordnet data 76.99% 51.21%
Trail 2, random wordnet data 77.32% 50.57%
Trail 3, random wordnet data 77.30% 50.06%
Landmark wordnet data 77.35% 52.09%
Table 6: Performance Comparison: Landmark vs. Random Data for WOS.
Similarly, for the results shown in Table 6, a significant performance improvement was observed for both landmark and random data in Domain Match and Area Match after applying WordNet augmentation. Notably, the landmark-WordNet data achieved the highest accuracy in both Domain Match (77.35%) and Area Match (52.09%), representing increases of 3.5% and 4.61%, respectively. The significant improvement in Area Match for the landmark-WordNet data is particularly noteworthy, as Area Match is a more granular metric that requires higher precision from the model. Without augmentation, the landmark data slightly underperformed compared to the random data in both metrics. Similar to the Reuters dataset, this may be because the landmark data selection focuses on capturing core themes within the dataset, which may not always align perfectly with the overall distribution of the majority of samples. In contrast, random data selection tends to better reflect the overall distribution of the dataset, especially when the sample size is small, making it more likely to cover the main categories.
8 Ablation: Test the performance of Augmentation Using Machine Learning Model
In addition to evaluating the improvements afforded by LLMs, we investigated enhancements achieved through traditional machine learning models. Specifically, we employed three machine learning algorithms: Gradient Boosting Machine (GBM), LightGBM, and Random Forest, all built upon BGE-m3 embeddings of the documents. We then applied an ensemble approach to combine the outputs of these traditional models when applied to the embedded data, to determine the final category assignments. For the categorization of Reuters Partial Match documents, we focused on the top two categories identified by our model, as fine-tuned LLMs predict an average of 1.14 categories per document, fewer than two categories.
Reuters WOS
Part Match All Match Domain Match Area Match
Landmark Data 78.54% 70.06% 66.97% 19.03%
WordNet Data 78.04% 68.03% 65.97% 24.26%
LLM Rewrite 77.15% 67.56% 66.74% 25.72%
RAG Augment 88.46% 79.65% 73.60% 39.46%
Combine Data 89.11% 79.61% 73.91% 40.60%
Table 7: Performance Comparison of the Machine Learning Ensemble Model with Various Data Augmentation Techniques.
As illustrated in Table 7, when working with traditional machine learning models, data augmentation methods that result in a high level of embedding similarity can actually degrade model performance. Conversely, the RAG augmentation method, which generates diverse variations of the data while preserving semantic coherence, significantly boosts performance. The RAG method achieves the highest accuracy for the Reuters All Match and WOS Domain datasets, underscoring its effectiveness in enhancing traditional machine learning models.